OpenStudy (anonymous):

What is the solution to the system of equations? x=3y+5 x-4y=3

rishavraj (rishavraj):

take variable to one side/.,.... x - 3y = 5 x - 4y = 3 now just subtract the two equation

jimthompson5910 (jim_thompson5910):

Alternative way x-4y=3 3y+5-4y=3 ... replace x with 3y+5 solve for y and use this solution for y to get the solution for x

OpenStudy (anonymous):

m the first equation by -1 -x+3y= -5 x-4y=3 then subtract to find y

OpenStudy (anonymous):

thanks got it:)